A 7kg cannon ball makes a head on collision with second bowling ball initially at rest. The 7kg ball stops and the second ball rolls away from the collision with a velocity that is 2/3 the speed of the original ball before the collision. What is the mass of the second ball?
1
Expert's answer
2018-03-15T17:14:08-0400
Using the law of momentum conservation in a collision, we get
m_1 v_1+m_2 v_2=m_1 v_1'+m_2 v_2'
So
7v_1+m_2 0=m_1 0+m_2 2/3 v_1
7=〖2/3 m〗_2
m_2=3/2×7=10.5 kg